The Quick Mart has two kinds of nuts. Pecans sell for $1.55 per pound and walnuts sell for $1.95 per pound. How many pounds of walnuts must be added to 15 pounds of pecans to make a mixture that sells for $1.75 per pound?
Answer by stanbon(75887)

Pecans sell for $1.55 per pound and walnuts sell for $1.95 per pound. How many pounds of walnuts must be added to 15 pounds of pecans to make a mixture that sells for $1.75 per pound?
-----------------
Pecan DATA:
Amt = 15 lbs ; Value is 1.55*15 = 23.25 dollars
------------
Walnut DATA:
Amt = x lbs ; value is 1.95x dollars
-------------
Mixture DATA:
Amt = 15+x lbs ; value = 1.75(15+x) = 26.25 + 1.75x dollars
--------------
EQUATION:
value + value = value
23.25 + 1.95x = 26.25+1.5x
0.45x = 3
x = 6 2/3 lbs (amount of walnuts tht must be added to make the mix)
=================
Cheers,
Stan H.